Suppose Intel stock has a beta of 1.52​, whereas Boeing stock has a beta of 0.76. If the​ risk-free interest rate is 4.4 % and the expected return of the market portfolio is 13.1 %​, according to the​ CAPM, a. What is the expected return of Intel​ stock? b. What is the expected return of Boeing​ stock? c. What is the beta of a portfolio that consists of 65 % Intel stock and 35 % Boeing​ stock? d. What is the expected return of a portfolio that consists of 65 % Intel stock and 35 % Boeing​ stock? (There are two ways to solve​ this.)
Answer #1

a. Expected return of Intel​ stock = Risk-free rate + Beta(Market return - Risk-free rate) = 0.044 + 1.52(0.131 - 0.044) = 0.176 or 17.6%

b. Expected return of Boeing​ stock = Risk-free rate + Beta(Market return - Risk-free rate) = 0.044 + 0.76(0.131 - 0.044) = 0.110 or 11.0%

c. Portfolio beta = (0.65 × 1.52) + (0.35 × 0.76) = 1.3

d. Expected return of a portfolio = (0.65 × 0.176) + (0.35 × 0.110) = 0.153 or 15.3%